Mr. Akbar provides you with the following information-(all the transactions are


separate and independent of each other)


 Started business with cash Rs150000


 Purchased goods for cash Rs 25000


 Sold goods to C on credit Rs 20000


 Paid salary for cash Rs15000


 Deposited cash into the bank account Rs100000Identify the accounts being affected in the monetary transaction and Identify the type of


accounts identified - real, personal or nominal


b. Discuss the rule of passing the journal entry applicable here and pass the journal entry


(Golden rule or transaction analysis, any of these rule/s can be taken as a base to justify


the answer)

1
Expert's answer
2021-11-17T08:57:02-0500

1.

1) started business with cash 150,000

Accounts involved -> capital A/C = Personal Account

Cash A/C = Real Account


* Capital Account is a personal Account as it is related to a person, and Cash Account is real Account as it is related to Assets.

2)

purchased goods far cash. RS 25,000

Accounts involved are,purchase A/C=Nominal,

Cash A/C=Real.

cash are assets therefore, are real Account

purchase is an expense so, Nominal A/C.

3)

Sold goods to C on credit Rs 20,000

Accounts involved are, sales A/C =nominal,

Debtor (i.e. C) A/C = personal.

Sales is an Income therefore nominal Account and debtor

(i.e. C) is related to person therefore, personal A/C.

4)

paid salary for cash Rs 15,000

Accounts involved are Salary A/C = Nominal,

Cash A/C = Real

Salary is an expense therefore Nominal A/C & Cash is

an Asset therefore real account


5)

Deposited cash into Bank A/C. 115.000

Cash A/C = Real Account

Bank A/C= personal account.

Real Account as it's asset and Bank account is personal Account as it relates to a person or organisation.



2.

1)

Rule for passing the Journal Entry as per golden Rule.

1) personal Account

Debit - Receiver

Credit - Giver

2)

Real Account

Debit - What comes in

credit - what goes out.

3) Nominal Account.

Debit - All Expenses & losses

Credit - All Income & profit.


JOURNAL ENTRIES

1)started business with cash 15,000.

cash A/C(Debit) - 15000

To capital A/C (credit) - 15,000


2) purchased goods for cash Rs. 25,000

purchase A/C - Debit 25,000

TO cash A/C (credit)- 25,000


3) Sold goods to C on credit RS 20,000

Debtor A/C or C's A/C (Debit) - 20,000

To sales A/C(credit) - 20,000


4) paid salary for cash. Rs 15,000.

Salary A/C (debit) 15,000

To cash A/C (credit) 15,000


5) Deposited cash into Bank

Bank A/C (Debit) 15,000

To cash A/C (Credit) 15,000